OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (cwrw238):

slope of the line will be 2 use y - y1 = m(x - x1) here x1 = 5 and y1 = 3 y - 3 = 2(x - 5) y - 3 = 2x - 10 y = 2x - 7